OpenStudy (jiteshmeghwal9):

\(\pi\) is ratio of circumference & diameter of a circle. For every circle its value is approximately same i.e; 3.14... But it is counted as an irrational number having specific value ??

OpenStudy (kropot72):

\pi is the best-known transcendental number. The information here could be of interest: https://en.wikipedia.org/wiki/Transcendental_number

OpenStudy (unklerhaukus):

3

OpenStudy (zzr0ck3r):

Irrational numbers have specific value.

OpenStudy (skullpatrol):

All real numbers have a specific value.

OpenStudy (skullpatrol):

$$\Huge \pi \approx \dfrac{22}{7}$$

OpenStudy (skullpatrol):

It is counted as irrational because it cannot be exactly expressed as a ratio of two integers

OpenStudy (phi):

*** π is ratio of circumference & diameter of a circle. For every circle its value is approximately same i.e; 3.14... *** I would reword that to say π is ratio of circumference & diameter of a circle. For every circle its value is exactly the same i.e; approximately 3.14...

OpenStudy (jiteshmeghwal9):

but still the accurate value of \(\pi\) is unknown . the value of \(\pi\) for any circle is about 22/7, not exactly 22/7 .

OpenStudy (mathstudent55):

It is not unknown. It is well know. It is \(\pi\). It just cannot be expressed as a fraction of integers, but it is still known.

OpenStudy (skullpatrol):

It has been calculated to 10 trillion decimal places and can be calculated to as many decimal places as you want. How does that make the accurate value unknown?

OpenStudy (zzr0ck3r):

This is the idea that something is there even though we can't express it in a "natural" way. Take a right triangle where the arms off of the 90 degree angle are length 1 (1cm, 1ft,1mile,1...) The other arm is sqrt{2}. This is also irrational but we can see it.

imqwerty (imqwerty):

For circles they just use pi as 3.14 for convenience cuz no one will sit multiplying all digits of pi andd if you mean to say that the practically measured ratio is 3.14 then it is not exactly 3.14 Checkout the proof for statement pi is irrational